OpenStudy (anonymous):

PLEASE HELP!!! A segment with endpoints F(4,9) and G(7,2) is divided by a point H such that FH and GH form a 1:3 ratio. Find the Y value for H.

OpenStudy (amistre64):

if we find G-F, and divide it by 3, .... then add F back to it, we should find it

OpenStudy (amistre64):

G(7,2) -F (4,9) --------- (3,-7)/3 = (1,-7/3) +F(4, 9 ) ----------- H=(5, this)

OpenStudy (amistre64):

guess i cut instead of copied ... (3,-7)/3 = (1,-7/3) ... thereofre etc ...

OpenStudy (mrnood):

@amistre64 I think you need to divide by 4 , not 3 - then add 1/4 or 3/4 to get the point H the ratio is 1:3 therefore the division is into 4 equal sections

OpenStudy (amistre64):

yes, divide by 4 since we want 4 parts

OpenStudy (mrnood):

but you sdaid divide by 3.....

OpenStudy (amistre64):

adding 1/4 back to F gets is where we want to be

OpenStudy (amistre64):

it was late, the asker wasnt around, prolly just rushed that in my head is all.